INFRARED PHOTODISSOCIATION SPECTROSCOPY OF SINGLY CHARGED TRANSITION METAL NI-TROGEN CLUSTERS

\underline{E.D. PILLAI}, T.D. JAEGER and M. A. DUNCAN, University of Georgia, Department of Chemistry, Athens, GA, 30602-2556.

Metal cation-nitrogen clusters, M^{+}(N_{2})_{n}, of the first row transition metals are produced in a laser vaporization pulsed nozzle cluster source, size selected and excited in the infrared with an infrared OPO/OPA. Infrared photodissociation occurs by elimination of whole N_{2} molecules and the photofragement yield versus wavelength produces IR resonanceenhanced photodissociation (REPD) spectra for the clusters. Vibrational bands seen in the 2100 to 2400 cm^{−}^{1} region correspond to the N-N stretch but are shifted from those of free N_{2}. Density functional theory calculations on the bond enegeries, geometries and frequency shifts for first row M^{+} (N_{2}) are also discussed in regard to periodic trends.